OpenStudy (anonymous):

integrate (1/(x^2+2x+2)^2) dx

OpenStudy (anonymous):

(x^2+2x+2+1-1)^2=((x+1)^2+1)^2 if you let u=x+1 you get 1/(u^2+1)^2 you should get tan^-1(u^-3) so your answer should be \[(1/-3)\tan^{-1}( (x+1)^-3)\]

OpenStudy (anonymous):

the back of the book is showing a different answer...

OpenStudy (anonymous):

whats it showing i didnt do this on scratch paper. does it involve tan_1 or an ln

OpenStudy (anonymous):

\[1/2(\tan^{-1} (x+1))+ 1/2((x+1)/(x^2+2x+2))+c\]

OpenStudy (cruffo):

complete the square of the denominator - \[x^2 + 2x + 2 = (x+1)^2 +1\] Then let \[x+1 = \tan(y)\] \[dx = \sec^2(y)dy\] back sub to find \[\int \frac{1}{(\tan^2(y)+1)^2}\sec^2(y) dy\] Simplify to get \[\int \frac{1}{sec^2(y)}dy\] Simplify some more \[\int \cos^2(y)\; dy\] Use power reduction \[\frac{1}{2}\int (1-\cos(2y)) dy\]

OpenStudy (anonymous):

i just punch this in a derivitive calc and it spit out some thing different than both of ours,

OpenStudy (anonymous):

hes right creative thinking right there i forgot to do the d(tanu)

OpenStudy (cruffo):

Integrating you get \[\frac{1}{2}y - \frac{1}{4}\sin(2y) +c\] probably want to write that as \[\frac{1}{2}y - \frac{1}{2}\sin(y)\cos(y) +c\] Now back to the x's, got to set up the triangle.

OpenStudy (cruffo):

\[\text{Since }\;\;x + 1 = \tan(y), \text{ then }\;\; y = \tan^{-1}(x+1)\] And \[\sin(y) = \frac{x+1}{\sqrt{(x+1)^2+1}}\] \[\cos(y) = \frac{1}{\sqrt{(x+1)^2+1}}\]

OpenStudy (cruffo):

In total, the answer looks like what avenged56 wrote above.

OpenStudy (anonymous):

yes i would agree

OpenStudy (anonymous):

ill follow through with it and figure out whats going on from there. thanks for the help